David opened his piggy bank and found 3 pennies, 5 nickels, and $2.95 in dimes and quarters. Find the greatest number of quarters he could have if he had an odd number of dimes.
Answer by stanbon(48558)

David opened his piggy bank and found 3 pennies, 5 nickels, and $2.95 in dimes and quarters. Find the greatest number of quarters he could have if he had an odd number of dimes.
----------
Let number of dimes be "x"; Value of these is 10x cents
Let number of quarters be "y"; Value of these is 25y
--------
EQUATION:
value + value = 295 cents
10x+25y=295
--------
Make a list of odd multiples of 10: 10,30,50,70...etc
Make a list of multiples of 25: 25,50,75,100...etc
------
Find the largest quarter multiple that when added to an odd multiple of 10
gives you 295: 295-10=285 no good
295-30=265 no good
295-50=245 no good
295-70=225 Ah, good because 9*25=295
-------
so, largest number of quarters would be 9
=========
Cheers,
Stan H.